HAROLD INVESTED $8000 FOR ONE YEAR, PART AT 11% SIMPLE INTREST AND PART AT 10% INTREST. IF HE EARNED A TOTAL OF $850 IN INTREST AT EACH RATE?
FOR 11% I GOT 5000
FOR 10% I GOT 3000
IS THIS Correct?

Answer by Boreal(15235)

x at 11%
8000-x at 10%
.11x+.10(8000-x)=850
.11x+800-.10x=850
.01x=50
x=$5000
8000-x=$3000
yes, you are correct.